OpenStudy (anonymous):

please help Create your own unique literal equation and solve for one of the variables. Show your work. Then, using complete sentences, explain how you solved for the variable you chose.

OpenStudy (paxpolaris):

write any equation with more-than-one variable.

OpenStudy (anonymous):

like 3x+5y=

OpenStudy (paxpolaris):

sure, \(3x+5y=1\) for example. then choose one of the variables and solve for it. let's say you choose to solve for x you want to rewrite the equation so you are just left with x on one side of the equal sign, and everything else on the other \[\Large x=\ \ \ \dots\]

OpenStudy (paxpolaris):

\[3x+5y=1\\ \implies 3x= 1-5y \\ \large \implies x={1-5y \over3}\]